# Comprehensive Evaluation of Prototype Neural Networks

### v1

@@ Abstract
Prototype models are an important method for explainable artificial intelligence (XAI) and interpretable machine learning. In this paper, we perform an in-depth analysis of a set of prominent prototype models including ProtoPNet, ProtoPool and PIPNet. For their assessment, we apply a comprehensive set of metrics. In addition to applying standard metrics from literature, we propose several new metrics to further complement the analysis of model interpretability. In our experimentation, we apply the set of prototype models on a diverse set of datasets including fine-grained classification, Non-IID settings and multi-label classification to further contrast the performance. Furthermore, we also provide our code as an open-source library, which facilitates simple application of the metrics itself, as well as extensibility – providing the option for easily adding new metrics and models. https://github.com/uos-sis/quanproto
@@
One possible solution to overcome some of these problems, has been the development of deep part-based prototype models, firstly presented in [15] and [16]. A key advantage is their intrinsic interpretability. Part-based prototype models focus on learning a meaningful embedding space using an often pre-trained black-box network as feature extractor. A key distinction from standard black-box models is the aim to extract features that concentrate on specific sections of the image. This division of the image in the embedding space allows the network to learn prototypes that represent only particular parts of the object. The classification layer then utilizes the presence of a prototype in the image to calculate the prediction. To provide an explanation for a prediction, the prototypes that have the highest evidence are projected back to the input space through a saliency method. This simple visual representation of important image parts typically allows even non-technical users to understand the model’s decision process.
@@ 2 Related Work
Below, we start with a brief summary of concept-based artificial intelligence, before describe architectures, limitations, and evaluation methods of deep learning-based part-prototype models. For a comprehensive survey on classical prototype models, we refer to Biehl et al. [24].
@@
The prototypical part network (ProtoPNet) from Chen et al. [16] is one of the pioneering networks that popularized the concept of part prototype networks in image classification. In addition to the models examined in this article, other networks have been developed that combine the prototype approach with many concepts from the ML field. Nauta et al. [32] varied the classification process, by replacing the fully connected layer at the end of the model through a decision tree (ProtoTree) [32]. ProtoPShare [33] uses an iterative post-processing technique to merge prototypes from a ProtoPNet model, effectively creating shared prototypes between classes and reducing the amount of needed prototypes. Xue et al. [34] adapted the idea of ProtoPNet to be used together with vision transformers as feature extractors. Prototype-based deep learning models have also been used for time series [35], graph data [36], or image segmentation tasks [37]. There is also work, which combines deep prototype models with an autoencoder architecture [38]. Other variants of the ProtoPNet architecture, include [39], [40], [41], [42].
@@
Part-based prototype models have been used e. g., to inspect power grids visually [43], in the classification of MRI scans to detect Alzheimer’s disease [44] or X-Ray images to classify Chest X-rays [45]. Carlone et al. [46] propose a prototype-based model to identify breast cancer, highlighting the acceptance of visual explanations.
@@
Besides the success of ProtoPNet and the improvement through architectural changes, there are also limitations and some critique regarding the way the prototypical networks explain their decisions. Xu-Darme et al. [47] criticize the way ProtoPNet and ProtoTree visualize the learned prototypes and propose to use different saliency methods to overcome this problem. Hoffmann et al. [48] point out that ProtoPNets struggle with adversarial examples or compression artifacts in the input. Nevertheless, Hoffmann et al. do not want to dispute part prototypical networks, but raise the awareness of these problems.
@@ 3 Datasets
For general object detection in a Non-IID setting, we used the NICO Dataset from He et al. [22]. This dataset includes 10 animal classes and 9 vehicle classes, though our focus is on the animal classes. Each class is further divided into 10 context categories representing various environments, attributes, and activities. The dataset contains 12,980 images of animals. Essentially, this Non-IID setting is a more realistic scenario, which is also new for part-based prototypical models. Hence, we want to apply this setting mainly through different background environments in the training and test sets to assess if the networks naturally promote object-focused prototypes. Since this cannot be achieved using the provided context categories (as few represent distinct environments), we use a simple image analysis inspired by Cheng et al. [54], based on the HSV color space [55], to split the dataset into training/validation and test sets. This approach effectively creates a Non-IID scenario with varying distributions across training and test data. Further details are given in the implementation.
@@
The last dataset is Animals with Attributes 2 (AWA2) introduced by Xian et al. [23]. This dataset is a benchmark for attribute-based classification, a related model domain to part-based prototype models. The dataset comprises 37,322 images across 50 animal classes, with binary and continuous class attributes. For evaluating the performance of part-based prototype methods in a multi-label classification setting, we focus on the given attributes. The AWA2 dataset includes 85 attributes that are used to describe a class. Since not all attributes describe visible features, we only use a subset of all attributes, i. e., those corresponding to the 49 visible attributes.
@@ 4 Prototype Networks
Because this work focuses on a comprehensive set of metrics. We kept our set of models small and selected three prominent networks based on the following criteria: class-specific vs. shared prototypes and explicit vs. indirect prototype representations. We chose ProtoPNet [16], the first popular prototype model, that uses class-specific explicit prototype vectors. We further include the ProtoPool [18] model as a representative for shared and explicit prototype learning. Finally, we chose PIPNet [19] as a recently introduced architecture for the shared and indirect case.
Each network uses a similar architecture visualized in Figure 1, the components are labelled with the letters (A)-(F) in the figure: Starting with a deep neural network (A) as backbone. This network extracts a feature map (B) from the input image. The important characteristic of a part-based prototype model is to calculate the presence of prototypes in the feature map. In the case of a model with specific prototypes, each prototype (C) is compared with the feature vectors in (B) resulting in a set of similarity maps (D). In the case of indirect prototypes, the feature map is further processed and directly interpreted as a collection of similarity maps (D) without an explicit comparison. The final classification (F) is then obtained based on the highest similarity score (E) in each map.
@@
The first network is ProtoPNet. It follows the general architecture shown in Figure 1 consisting of a convolutional neural network as a backbone followed by two additional convolutional layers to adjust the output dimension . This results in a feature map with individual vectors . The feature map is then processed by a prototype layer, that holds a set of prototypes with , such that each prototype has the same dimensions as one feature vector. A pre-determined number of prototypes is allocated for each class . The prototype layer computes a set of similarity maps based on the squared L2 distance between a prototype and a feature vector as follows:
@@
ProtoPNet uses a multistep training process, separating the training of the last layer from the rest of the network. At the beginning, the last layer is initialized using for weights connecting a prototype to its corresponding class and otherwise. This ensures the learning of class specific prototypes. The first training step optimizes the backbone and prototype layer to learn a meaningful latent space, clustering prototypes from the same class and separating them from prototypes of other classes. This is done using a linear combination of a classification loss and two novel losses, the cluster loss and separation loss . In the second step, each prototype is first projected onto the nearest feature vector from its assigned class. The last layer is then fine-tuned to improve classification performance, using only the classification loss and an L1 regularization term to keep the positive reasoning characteristic. Chen et al. [16] observed that some prototypes focus on the background. These prototypes contradict the goal of the model to learn class-specific object parts. To address this, the authors propose a pruning step after training, which removes such prototypes. We use both the original and pruned version in the evaluation to analyze the effectiveness of this approach.
@@
The ProtoPool [18] model by Rymarczyk et al., addresses several limitations of ProtoPNet. It introduces a new similarity function and an automatic, fully differentiable assignment of prototypes to classes, that enables the network to learn shared prototypes. The model itself follows the architecture of ProtoPNet consisting of a CNN backbone with additional convolutional layer to adjust the output shape, a prototype layer, and a fully connected layer. The prototype layer contains a set of trainable prototypes with , similar to ProtoPNet. A novel part is the assignment matrix with prototype slots for each class, where each slot is implemented as a distribution over all available prototypes. The similarity maps are computed using the same equation 1 as in ProtoPNet. To enhance prototype locality, the authors introduce a focal similarity function, which computes the difference between the maximum similarity and the average similarity as follows:
@@
PIPNet follows a multistep training process. In the first step, only the backbone is optimized using the alignment and tanh loss to learn prototypes representing general, class-independent features. In the second step, the entire network is optimized using all loss terms, including classification loss. To promote sparsity during training, ensuring that each class uses a minimal number of prototypes for classification, the output scores are calculated as follows:
@@
| O. Completeness | [57] | [57] | VLC [57] | VAC | |||
| Continuity | CAC | CRC | |||||
| Contrastivity | [58] | [58] | Entropy | ||||
| C. Complexity | O. Overlap [58] | B. Overlap | IORD [58]* | Cons. [49]* | |||
| Compactness | Global Size [19] | Sparsity [19] | NPR | Local Size [19] |
Multi-Label Optimization
The AWA2 dataset has multiple labels per image. To adapt the training objectives for this setting, we switch from a cross-entropy loss to a multi-label margin loss as the classification loss. Other common loss functions for multi-label classification include binary cross-entropy loss and multi-label soft margin loss. However, these losses tend to learn negative weights in the classification layer for labels not present in the image, which conflicts with the L1 regularization used to promote positive reasoning in the final layer. Other class specific loss functions that need to be adapted are the cluster and separation loss used in ProtoPNet and ProtoPool. We calculate the mean over the individual cluster and separation losses of each label. We argue that this simple extension remains true to the idea of these losses described in ProtoPNet [16].
@@ 4.1 Prototype Visualization
The post-hoc visualization method is not bound to the individual models, allowing us to select the most suitable method for the networks. The original implementation of ProtoPNet employs a simple model-agnostic upscaling method, which was also used by ProtoPool and PIPNet. This method upscales the similarity map of a prototype to the size of the input using a cubic interpolation to visualize the similarity pattern of the prototype within the image. A key disadvantage is that it relies solely on the similarity map and assumes the preservation of spatial information in the feature extraction process. There are other model-specific methods that can better visualize the feature extraction process of the model. Based on these arguments and the findings from Xu-Darme et al.[59], which compared various visualization methods, we selected the Prototypical Relevance Propagation PRP method from Gautam et al.[60] for visualizing the prototypes.
@@ 5 Evaluation Metrics
In this section, we discuss the techniques and metrics used to evaluate the part-based prototype models. Following the Co-12 properties from Nauta et al. [17] as a guideline, we choose metrics to evaluate the output-completeness, continuity, contrastivity, covariate complexity, and compactness. We use a total of 22 metrics, including 13 metrics that are novel to this domain, an overview of the metrics categorized by the evaluation techniques is shown in Table 1. We only use the top-5 most activated prototypes of a sample image to evaluate prototype quality-related metrics. This approach is aligned with the actual explanation scheme, where only a few prototypes are visualized to explain a decision to the user. To evaluate the general performance of the networks, we use accuracy, top-3 accuracy, and F1 score.

### v2

@@ Abstract
Prototype models are an important method for explainable artificial intelligence (XAI) and interpretable machine learning. In this paper, we perform an in-depth analysis of a set of prominent prototype models including ProtoPNet, ProtoPool and PIPNet. For their assessment, we apply a comprehensive set of metrics. In addition to applying standard metrics from literature, we propose several new metrics to further complement the analysis of model interpretability. In our experimentation, we apply the set of prototype models on a diverse set of datasets including fine-grained classification, Non-IID settings and multi-label classification to further contrast the performance. Furthermore, we also provide our code as an open-source library (https://github.com/uos-sis/quanproto), which facilitates simple application of the metrics itself, as well as extensibility – providing the option for easily adding new metrics and models.
@@
One possible solution to overcome some of these problems, has been the development of deep part-based prototype models, firstly presented in [15] and [16]. A key advantage is their intrinsic interpretability. Part-based prototype models focus on learning a meaningful embedding space using an often pre-trained black-box network as feature extractor. A key distinction from standard black-box models is the aim to extract features that concentrate on specific sections of the image. This division of the image in the embedding space allows the network to learn prototypes that represent only particular parts of the object. The classification layer then utilizes the presence of a prototype in the image to calculate the prediction. To provide an explanation for a prediction, the prototypes that have the highest evidence are projected back to the input space through a saliency method.
However, users have different experiences with model explanations. Confirmation bias and subjective understanding make it hard to guess the prediction correctly [17] [18]. This shows that these models are not yet reliable enough to use for AI-assisted decision-making. This encouraged us to survey existing assessments of explainability using a well-founded categorization.
@@ 2 Related Work
This work addresses a broad spectrum of subdomains that are relevant to prototype-based networks. Central to this line of research is the idea of learning human-interpretable concepts, a property particularly important in high-stakes applications. While prototype-based models represent a widely adopted approach in this domain, they also exhibit notable limitations. These shortcomings highlight the need of practical and comprehensive analysis tools.
@@
The prototypical part network (ProtoPNet) from Chen et al. [16] is one of the pioneering networks that popularized the concept of part prototype networks in image classification. In addition to the models examined in this article, other networks have been developed that combine the prototype approach with many concepts from the ML field. Nauta et al. [33] varied the classification process, by replacing the fully connected layer at the end of the model through a decision tree (ProtoTree) [33]. ProtoPShare [34] uses an iterative post-processing technique to merge prototypes from a ProtoPNet model, effectively creating shared prototypes between classes and reducing the amount of needed prototypes. Xue et al. [35] adapted the idea of ProtoPNet to be used together with vision transformers as feature extractors. Prototype-based deep learning models have also been used for time series [36], graph data [37], or image segmentation tasks [38]. There is also work, which combines deep prototype models with an autoencoder architecture [39]. Other variants of the ProtoPNet architecture, include [40], [41], [42], [43]. Li et al. [44] provide a comprehensive overview of the design choices and prototype formulations that are used in the literature analysing prototype quality relations.
@@
Part-based prototype models have been used in several industrial application settings e. g., to inspect power grids visually [45], in the classification of MRI scans to detect Alzheimer’s disease [46] or X-Ray images to classify Chest X-rays [47]. Carlone et al. [48] propose a prototype-based model to identify breast cancer, highlighting the acceptance of visual explanations.
@@
Besides the success of ProtoPNet and the improvement through architectural changes, there are also limitations and some critique regarding prototypical networks. Elhadri et al. [49] survey general quality issues regarding the interpretability of learned prototypes in these models. Xu-Darme et al. [50] criticize the way ProtoPNet and ProtoTree visualize the learned prototypes and propose to use different saliency methods to overcome this problem. Hoffmann et al. [51] point out that ProtoPNets struggle with adversarial examples or compression artifacts in the input. Nevertheless, Hoffmann et al. do not want to dispute part prototypical networks, but raise the awareness of these problems. Saralajew et al. [52] contribute to this discussion by introducing a method for fast and provable adversarial robustness certification for Nearest Prototype Classifiers (NPCs).
@@ 3 Datasets
For general object detection in a Non-IID setting, we used the NICO Dataset from He et al. [24]. This dataset includes 10 animal classes and 9 vehicle classes, though our focus is on the animal classes. Each class is further divided into 10 context categories representing various environments, attributes, and activities. The dataset contains 12,980 images of animals. Essentially, this Non-IID setting is a more realistic scenario, which is also new for part-based prototypical models. Hence, we want to apply this setting mainly through different background environments in the training and test sets to assess if the networks naturally promote object-focused prototypes. Since this cannot be achieved using the provided context categories (as few represent distinct environments), we use a simple image analysis inspired by Cheng et al. [58], based on the HSV colour space [59], to split the dataset into training/validation and test sets. This approach effectively creates a Non-IID scenario with varying distributions across training and test data. Further details are given in the implementation.
@@
The last dataset is Animals with Attributes 2 (AWA2) introduced by Xian et al. [25]. This dataset is a benchmark for attribute-based classification, a related model domain to part-based prototype models. The dataset comprises 37,322 images across 50 animal classes, with binary and continuous class attributes. For evaluating the performance of part-based prototype methods in a multi-label classification setting, we focus on the given attributes. The AWA2 dataset includes 85 attributes that are used to describe a class. Since not all attributes describe visible features, we only use a subset, i. e., those corresponding to the 49 visible attributes.
@@ 4 Prototype Networks
Because this work focuses on a comprehensive set of metrics, we kept our set of models small and selected three prominent networks based on the following criteria: class-specific vs. shared prototypes and explicit vs. indirect prototype representations. We chose ProtoPNet [16], the first popular prototype model, that uses class-specific explicit prototype vectors. We further include the ProtoPool [20] model as a representative for shared and explicit prototype learning. Finally, we chose PIPNet [21] as a recently introduced architecture for the shared and indirect case. To the best of our knowledge, there is currently no model for the class-specific indirect case. Other models with different classification approaches, such as hierarchical classifications or other backbones such as transformers, are further important studies that should be investigated in connection with the interpretability of prototypes. However, due to a lack of implementations of visual methods and complex adjustments for the multi-label scenario, we decided to start with models that can be easily adapted to our training and evaluation setups.
Each network uses a similar architecture visualized in Figure 1, the components are labelled with the letters (A)-(G) in the figure: Starting with a deep neural network (A) as backbone. This network extracts a feature map (B) from the input image. The important characteristic of a part-based prototype model is to calculate the presence of prototypes in the feature map. In the case of a model with specific prototypes, each prototype (C) is compared with the feature vectors in (B) resulting in a set of similarity maps (D). In the case of indirect prototypes, the feature map is further processed and directly interpreted as a collection of similarity maps (D) without an explicit comparison. The final classification (F) is then obtained based on the highest similarity score (E) in each map. Another important characteristic of these models is the visualization of such prototypes (G). In our case, this is done with the Prototypical Relevance Propagation (PRP) method that propagates the similarity score back to the input.
@@
The first network is ProtoPNet. It follows the general architecture shown in Figure 1 consisting of a convolutional neural network as a backbone followed by two additional convolutional layers to adjust the output dimension . This results in a feature map with individual vectors . The feature map is then processed by a prototype layer, that holds a set of prototypes with , such that each prototype has the same dimensions as one feature vector. A pre-determined number of prototypes is allocated for each class . The prototype layer computes a set of similarity maps based on the squared L2 distance between a prototype and a feature vector.
@@
ProtoPNet uses a multistep training process, separating the training of the last layer from the rest of the network. At the beginning, the last layer is initialized using for weights connecting a prototype to its corresponding class and otherwise. This ensures the learning of class specific prototypes. The first training step optimizes the backbone and prototype layer to learn a meaningful latent space, clustering prototypes from the same class and separating them from prototypes of other classes. This is done using a linear combination of a classification loss and two novel losses, the cluster loss and separation loss . In the second step, each prototype is first projected onto the nearest feature vector from its assigned class. The last layer is then fine-tuned to improve classification performance, using only the classification loss and an L1 regularization term to keep the positive reasoning characteristic. Chen et al. [16] observed that some prototypes focus on the background. These prototypes contradict the goal of the model to learn class-specific object parts. To address this, the authors propose a pruning step after training, which removes such prototypes. We use both the original and pruned version in the evaluation to analyse the effectiveness of this approach.
Our experiments include a multi-label dataset. To adapt the training objectives for this setting, we switch from a cross-entropy loss to a multi-label margin loss. Other common loss functions for multi-label classification include binary cross-entropy loss and multi-label soft margin loss. However, these losses tend to learn negative weights in the classification layer for labels not present in the image, which conflicts with the L1 regularization used to promote positive reasoning in the classification layer. The multi-label adaption for the cluster and separation loss is done by calculating the mean cluster and separation loss across all classes , i.e., the set of classes assigned to sample .
We argue that this is a fitting adaption that does not contradict with the goal described in the original study.
@@
The ProtoPool [20] model by Rymarczyk et al., addresses several limitations of ProtoPNet. It introduces a new similarity function and an automatic, fully differentiable assignment of prototypes to classes, that enables the network to learn shared prototypes. The model itself follows the architecture of ProtoPNet consisting of a CNN backbone with additional convolutional layer to adjust the output shape, a prototype layer, and a fully connected layer. The prototype layer contains a set of trainable prototypes with , similar to ProtoPNet. A novel part is the assignment matrix with prototype slots for each class, where each slot is implemented as a distribution over all available prototypes. The similarity maps are computed using the same equation 1 as in ProtoPNet. To enhance prototype locality, the authors introduce a focal similarity function, which computes the difference between the maximum similarity and the average similarity.
@@
The adaptation for multi-label training objectives follows the same approach as the ProtoPNet model. This involves switching from cross entropy loss to a multi-label margin loss and using the mean cluster and separation loss across all labels as illustrated in equation 2 and 3. No adaptation is necessary for the orthogonal loss , as it computes cosine similarities between each slot without utilizing class information.
@@
PIPNet follows a multistep training process. In the first step, only the backbone is optimized using the alignment and tanh loss to learn prototypes representing general, class-independent features. In the second step, the entire network is optimized using all loss terms, including classification loss. To promote sparsity during training, ensuring that each class uses a minimal number of prototypes for classification, the output scores are further modified.
@@
Since PIPNet’s loss functions are independent of class assignments, they can be used directly for multi-label classification. The only modification required is that the classification loss be calculated using the multi label margin loss rather than the cross entropy loss.
@@ 4.1 Prototype Visualization
The post-hoc visualization method is not bound to the individual models, allowing us to select the most suitable method for the networks. The original implementation of ProtoPNet employs a simple model-agnostic upscaling method, which was also used by ProtoPool and PIPNet. This method upscales the similarity map of a prototype to the size of the input using a cubic interpolation to visualize the similarity pattern of the prototype within the image. A key disadvantage is that it relies solely on the similarity map and assumes the preservation of spatial information in the feature extraction process. There are other model-specific methods that can better visualize the feature extraction process of the model. Based on these arguments and the findings from Xu-Darme et al.[61], which compared various visualization methods, we selected the Prototypical Relevance Propagation (PRP) method from Gautam et al.[62] for visualizing the prototypes (see Figure 1).
@@ 5 Evaluation Metrics
In this section, we discuss the techniques and metrics used to evaluate the part-based prototype models. Following the Co-12 properties from Nauta et al. [19] as a guideline, we choose metrics to evaluate the output-completeness, continuity, contrastivity, covariate complexity, and compactness.
Output Completeness: Evaluates whether the visualization covers all pixels relevant to a prototype by perturbing unhighlighted pixels and measuring the induced changes in the model and explanations. Measurement of maximum similarity location shift after perturbation. Large shifts indicate poor alignment between the visualization location and the nearest feature location in the feature map. [63] Quantifies the change in the maximum similarity value. A high change indicates that the relevant image region of the nearest feature is not fully captured by the visualization Measures the spatial similarity distribution shift across the feature map. Large deviations indicate poor alignment between visualization and spatial feature extraction. Quantifies changes in overall similarity magnitude irrespective of location. High values suggest that the visualization did not cover the image/object characteristics similar to the prototype. VLC [63] Measures displacement of the visualization’s highlighted region between original and perturbed images. Large displacements imply a strong impact of the induced spatial feature extraction changes on the visualization method VAC Quantifies changes in explanation activation between original and perturbed images. High values reveal a strong impact of the induced information extraction change on the visualization. Continuity: Assesses prototype stability under image-wide, low-level perturbations, indicating whether the model has learned robust, high-level semantic features. Measures the shift in the maximum-similarity location after global low-level perturbations. Large shifts imply a misalignment to the feature map interpretation as spatial aligned high level object part representations. [63] Quantifies changes in the maximum similarity value under global perturbations. High values indicate the model learned unstable prototypes representing low level image characteristics. Measures changes in the overall spatial similarity pattern after low-level perturbations. Large changes indicate spatial instability in the feature extraction process. [63]* Reports rank changes of the prototype score relative to other prototypes under perturbations. Large rank shifts indicate prototypes are influenced differently, indicating different degrees of semantic representation. Quantifies changes in overall similarity magnitudes after low-level perturbations. High values indicate large movement in the embedding space showing features represent low level image characteristics. CAC% Measures changes in output logits induced by low-level perturbations. Large variations indicate that nuisance characteristics substantially affect the classification process. CRC Reports rank changes of the predicted class under perturbations. Large rank drops indicate limited class-level robustness and reliance on such non-semantic information. Contrastivity: Assesses whether different prototypes capture distinct object properties, promoting robust classification and informative explanations. Measures the location distance between prototype similarity maxima in a sample image. Larger distances indicate that prototypes focus on distinct object parts. Measures spatial separation of prototype similarity pattern in a sample image. Higher values denote disentangled unique feature vectors resulting in focused prototype activations. Computes average pairwise distance between prototypes within a class. Larger distances indicate prototypes represent distinct object parts. Computes average distance between feature vectors closest to a prototype. Large deviations to the average prototype distance indicate a distorted class cluster distribution. [64] Measures average distance between prototypes across classes. Larger distances indicate well-separated class representations in the embedding space. [64] Measures average distance between feature vectors closest to a prototype across classes. Small values imply minimal class distinct information preserved in the feature extraction process. Entropy Computes the entropy of the similarity scores from a prototype across the test set. Lower entropy indicates better embedding space separation with only intended class features near the prototype. Covariate Complexity: Assesses the semantic focus of prototypes by quantifying overlap between prototype visualizations, object masks, and annotated parts. O. Overlap [64] Measures object mask coverage of visualizations. Lower values indicate part-level focus rather than broad object focus. B. Overlap Measures the background overlap of visualizations. Low overlap indicates minimal reliance on non-object context. IORD [64]* Measures the difference between average activations inside versus outside the object region. Higher scores indicate object-centric explanations rather than background reliance. Consistency [53]* Measures agreement between a prototype’s visual focus and annotated object parts across images. Higher consistency indicates stable and interpretable connection to semantic object parts. Compactness: Assesses model size and interpretability of the classification process. Global Size [21] Counts the number of active prototypes used by the model. Smaller values increase interpretability of the classification process by reducing its overall size. Sparsity [21] Measures sparsity of the classification layer. Higher sparsity indicates reliance on fewer prototypes per class, increasing interpretability. NPR Computes the ratio of negative to positive classification weights. Lower ratios improve interpretability by aligning decisions with part-based evidence and not absence. Local Size [21] Counts prototypes near the extracted features. Smaller values increase interpretability, meaning only a small number of prototypes are used for classification.
We use a total of 22 metrics, including 13 metrics that are novel to this domain, an overview of the metrics categorized by the evaluation techniques is shown in Table 1. The coverage of the model components by the metrics is shown in Figure 1, with additional information if the metrics focus on spatial, activation, or other features.
We only use the top-5 most activated prototypes of a sample image to evaluate prototype quality-related metrics. This approach is aligned with the actual explanation scheme, where only a few prototypes are visualized to explain a decision to the user. To evaluate the general performance of the networks, we use accuracy, top-3 accuracy, and F1 score.